The distribution of grades in an introductory finance class is normally distributed, with an expected grade of 68. If the standard deviation of grades is 15, in what range would you expect 68.26 percent of the grades to fall? (Round answers to 2 decimal places, e.g. 15.25. Hint: Think in terms of what the expected highest and lowest scores would be for 68.26% of the students taking the exam.)

68.26 percent of the grades should fall beween the ranges (lower end)?? enter the lower end of the range and (upper end)?? enter the upper end of the range.
Answer #1

The distribution of grades in an introductory finance class is normally distributed, with

expected grade u = 68

standard deviation of grades sd = 15

for 68.26% normally distributed, Z = +-1

So, range is calculated using (u-Z*sd, u+Z*sd)

Lower limit = 68 - 1*15 = 53

upper limit = 68 + 1*15 = 83

68.26 percent of the grades should fall beween the ranges 53 and 83